Profitability ratios: Cisco Systems has total assets of $35.594 billion, total debt of $9.678 billion, and net sales of $22.045 billion. Their net profit margin for the year was 20 percent, while the operating profit margin was 30 percent. What are Cisco’s net income, EBIT ROA, ROA, and ROE?
Answer #1

Total assets- $35.594 billion Total debt= 9.678 billion Net sales $22.045 billion Net profit margin-20% Operating profit marg
Answer #2

net income,

22.045 billion x 20% = 4.409 billion

EBIT

22.045 billion x 30% = 6.614 billion

ROA,

4.409 billion / 35.594 billion = 12.39%

ROE

35.594 - 9.678 = 25.916 billion total equity
4.409 / 25.916 = 17.01% RO
Answer #3

SOLUTION :



NP (net income)

= NPM * Sales 

= 0.20 * 22.045 

= 4.409 billion dollars. (ANSWER).



OP = EBIT 

= OPM * Sales 

= 0.30 * 22.045 

= 6.6135 billion dollars 


EBIT ROA 

= EBIT / Assets

= 6.6135/35.594

= 0.1858

= 18.58 % (ANSWER)


ROA

= NP / Assets

= 4.409 / 35.594

= 0.1239

= 12.39% (ANSWER)


Equity = Assets - Debt = 35.594 - 9.678 = 25.916 billion dollars.


ROE 

= NP / Equity

= 4.409 / 25.916

= 0.1701

= 17.01 % (ANSWER).



answered by: Tulsiram Garg